### v3

@@ 5.1 The linear case
Note that the randomness in (68) appears through $\alpha\in L_{\mathbb{F}}^{\infty}(\Omega;L^{\infty}(0,T;L^{\infty}(\mathcal{O})))$ and $q_{0}\in L^{p^{\prime}}(\Omega;L^{p^{\prime}}(\mathcal{O}))$. For $\omega\in\Omega$ fixed, we have from [FCZ00, Proposition 3.2]
@@
$\left\lVert q(T)\right\rVert_{L^{p^{\prime}}(\Omega;L^{p^{\prime}}(\mathcal{O}))}\leq C\left\lVert q\right\rVert_{L^{p^{\prime}}(\Omega;L^{p^{\prime}}((0,T)\times\mathcal{O}_{0}))}.$ (70)
@@
Now, we will follow an argument given by [TZ09], adapted to the $L^{p^{\prime}}$-setting. For any $y_{T}\in L^{p}(\Omega;L^{p}(\mathcal{O}))$, the goal is to find a control $h\in L^{p}(\Omega;L^{p}((0,T)\times\mathcal{O}_{0}))$ such that the solution $y$ of (67) satisfies $y(0)=0$.
We introduce the following linear subspace of $L_{\mathcal{F}}^{p^{\prime}}(0,T;L^{p^{\prime}}(\mathcal{O}_{0}))$,
@@
The observability estimates (70) leads to the fact that $\mathcal{L}$ is a bounded linear functional on $\mathcal{X}$ because
$\left|\mathbb{E}\int_{\mathcal{O}}y_{T}q(T)\,\textnormal{d}x\right|\leq\left\lVert y_{T}\right\rVert_{L^{p}(\Omega;L^{p}(\mathcal{O}))}\left\lVert q(T)\right\rVert_{L^{p^{\prime}}(\Omega;L^{p^{\prime}}(\mathcal{O}))}\leq C\left\lVert y_{T}\right\rVert_{L^{p}(\Omega;L^{p}(\mathcal{O}))}\left\lVert q\right\rVert_{L^{p^{\prime}}(\Omega;L^{p^{\prime}}((0,T)\times\mathcal{O}_{0}))}.$ (71)
By the Hahn-Banach theorem, $\mathcal{L}$ can be extended to a bounded linear functional on $L_{\mathcal{F}}^{p^{\prime}}(0,T;L^{p^{\prime}}(\mathcal{O}_{0}))$ with
$\|\mathcal{L}\|_{\mathcal{L}(L_{\mathcal{F}}^{p^{\prime}}(0,T;L^{p^{\prime}}(\mathcal{O}_{0})),{\mathbb{R}})}\leq C\left\lVert y_{T}\right\rVert_{L^{p}(\Omega;L^{p}(\mathcal{O}))},$ (72)
@@
$(L_{\mathcal{F}}^{p^{\prime}}(\Omega;L^{p^{\prime}}(0,T;L^{p^{\prime}}(\mathcal{O}_{0})))^{\prime}=L_{\mathcal{F}}^{p}(\Omega;L^{p}(0,T;L^{p}(\mathcal{O}_{0}))).$
This means that there exists $h\in L_{\mathcal{F}}^{p}(\Omega;L^{p}(0,T;L^{p}(\mathcal{O}_{0})))$ such that
$\mathcal{L}(q)=\mathbb{E}\int_{0}^{T}\int_{\mathcal{O}_{0}}hq\,\textnormal{d}x\textnormal{d}t\qquad\forall q\in L^{p^{\prime}}(\Omega;L^{p^{\prime}}(0,T;L^{p^{\prime}}(\mathcal{O}_{0}))),$ (73)
@@
$\mathcal{L}(q)=-\mathbb{E}\int_{\mathcal{O}}y_{T}q(T)\,\textnormal{d}x=\mathbb{E}\int_{0}^{T}\int_{\mathcal{O}_{0}}hq\,\textnormal{d}x\textnormal{d}t\qquad\forall q_{0}\in L^{p^{\prime}}(\Omega;L^{p^{\prime}}(\mathcal{O})).$ (74)
@@
$\mathbb{E}\int_{\mathcal{O}}y(0)q_{0}\,\textnormal{d}x=0\qquad\forall q_{0}\in L^{p^{\prime}}(\Omega;L^{p^{\prime}}(\mathcal{O})).$
@@
$\|h\|_{L_{\mathcal{F}}^{p}(\Omega;L^{p}(0,T;L^{p}(\mathcal{O}_{0})))}\leq C\left\lVert y_{T}\right\rVert_{L^{p}(\Omega;L^{p}(\mathcal{O}))},$
@@
Now we deal with the more subtle case $p=+\infty$. The difficulty is that one cannot use [LYZ12, Corollary 2.3] because the so-called Radon-Nikodym property is not satisfied by $L^{\infty}$, in particular $(L_{\mathcal{F}}^{1}(\Omega;L^{1}(0,T;L^{1}(\mathcal{O}_{0})))^{\prime}$ is not equal to $L_{\mathcal{F}}^{\infty}(\Omega;L^{\infty}(0,T;L^{\infty}(\mathcal{O}_{0})))$. We start from the $L^{1}$-observability estimate for solution $q$ to (68)
$\left\lVert q(T)\right\rVert_{L^{1}(\Omega;L^{1}(\mathcal{O}))}\leq C\left\lVert q\right\rVert_{L^{1}(\Omega;L^{1}((0,T)\times\mathcal{O}_{0}))},$ (75)
@@
$\|\mathcal{L}\|_{\mathcal{L}(L_{\mathcal{F}}^{1}(0,T;L^{1}(\mathcal{O}_{0})),{\mathbb{R}})}\leq C\left\lVert y_{T}\right\rVert_{L^{\infty}(\Omega;L^{\infty}(\mathcal{O}))},$ (76)
where $C$ is the same as in (75). In particular, $\mathcal{L}$ is a bounded linear functional on $L_{\mathcal{F}}^{2}(0,T;L^{2}(\mathcal{O}_{0}))$ because $L_{\mathcal{F}}^{2}(0,T;L^{2}(\mathcal{O}_{0}))\hookrightarrow L_{\mathcal{F}}^{1}(0,T;L^{1}(\mathcal{O}_{0}))$, therefore we can use the following fact coming from [LYZ12, Corollary 2.3],
$(L_{\mathcal{F}}^{2}(\Omega;L^{2}(0,T;L^{2}(\mathcal{O}_{0})))^{\prime}=L_{\mathcal{F}}^{2}(\Omega;L^{2}(0,T;L^{2}(\mathcal{O}_{0}))).$
This means that there exists $h\in L_{\mathcal{F}}^{2}(\Omega;L^{2}(0,T;L^{2}(\mathcal{O}_{0})))$ such that
$\mathcal{L}(q)=\mathbb{E}\int_{0}^{T}\int_{\mathcal{O}_{0}}hqdxdt\qquad\forall q\in L^{2}(\Omega;L^{2}(0,T;L^{2}(\mathcal{O}_{0}))),$ (77)
@@
$\mathcal{L}(q)=-\mathbb{E}\int_{\mathcal{O}}y_{T}q(T)dx=\mathbb{E}\int_{0}^{T}\int_{\mathcal{O}_{0}}hqdxdt\qquad\forall q_{0}\in L^{2}(\Omega;L^{2}(\mathcal{O})).$
@@
Our goal is then to prove that $h\in L_{\mathcal{F}}^{\infty}(\Omega;L^{\infty}(0,T;L^{\infty}(\mathcal{O}_{0})))$ and
$\|h\|_{L_{\mathcal{F}}^{\infty}(\Omega;L^{\infty}(0,T;L^{\infty}(\mathcal{O}_{0})))}\leq 2C\|y_{T}\|_{L^{\infty}(\Omega;L^{\infty}(\mathcal{O}))}.$ (78)
@@
$\left|\mathbb{E}\int_{0}^{T}\int_{\mathcal{O}_{0}}hqdxdt\right|\leq C\|q\|_{L^{1}(\Omega;L^{1}(0,T;L^{1}(\mathcal{O}))}\|y_{T}\|_{L^{\infty}(\Omega;L^{\infty}(\mathcal{O}))}\quad\forall q\in L^{2}(\Omega;L^{2}(0,T;L^{2}(\mathcal{O}_{0}))).$ (79)
@@
$q=\text{sign}(h)1_{|h|>2C\|y_{T}\|_{L^{\infty}(\Omega;L^{\infty}(\mathcal{O}))}}1_{\mathcal{O}_{0}},$
@@
$2C\|y_{T}\|_{L^{\infty}(\Omega;L^{\infty}(\mathcal{O}))}\mu_{\Omega,[0,T],\mathcal{O}_{0}}(|h|>2C\|y_{T}\|_{L^{\infty}(\Omega;L^{\infty}(\mathcal{O}))}),$
@@
$C\|y_{T}\|_{L^{\infty}(\Omega;L^{\infty}(\mathcal{O}))}\mu_{\Omega,[0,T],\mathcal{O}_{0}}(|h|>2C\|y_{T}\|_{L^{\infty}(\Omega;L^{\infty}(\mathcal{O}))}),$
@@
$2C\|y_{T}\|_{L^{\infty}(\Omega;L^{\infty}(\mathcal{O}))}\mu_{\Omega,[0,T],\mathcal{O}_{0}}(|h|>2C\|y_{T}\|_{L^{\infty}(\Omega;L^{\infty}(\mathcal{O}))})\\ \leq C\|y_{T}\|_{L^{\infty}(\Omega;L^{\infty}(\mathcal{O}))}\mu_{\Omega,[0,T],\mathcal{O}_{0}}(|h|>2C\|y_{T}\|_{L^{\infty}(\Omega;L^{\infty}(\mathcal{O}))}).$
@@
$\mu_{\Omega,[0,T],\mathcal{O}_{0}}(|h|>2C\|y_{T}\|_{L^{\infty}(\Omega;L^{\infty}(\mathcal{O}))})=0.$
@@
$|h|\leq 2C\|y_{T}\|_{L^{\infty}(\Omega;L^{\infty}(\mathcal{O}))}\quad\text{a.e. }(t,x)\in Q_{T},\ \text{a.s.}$
@@ 5.2 The nonlinear control problem
The proof of Theorem˜1.11 relies on the so-called source term method, originally developed in [LTT13] for the deterministic setting and later extended to SPDEs in [HSLBP22]. While [HSLBP22] already discusses this method for BSPDEs, here we provide a more comprehensive treatment and extend the analysis to our setting.
@@
where $C$ is the constant provided by Proposition˜1.9 (with $p=+\infty$). We then introduce the function
$\forall t\in(0,T],\quad\gamma(t)=e^{\frac{M}{T-t}},$ (82)
which blows up as $t\to T^{-}$. For some parameters $Q\in(1,\sqrt{2})$ and $R>Q^{2}/(2-Q^{2})$, we also define
@@
The proof follows the same steps of [HSLBP22, Proposition 2.5] with some adaptations to the backward case and the $L^{\infty}$-setting (cf. [LB20]).
@@
  • In the proof of Theorem˜1.11, we employed a multi-valued fixed point theorem, in contrast to the standard Banach fixed point argument used in Theorem˜1.7. The main reason for this is that Proposition˜5.1 guarantees only the existence of a control that steers the solution of (80) to zero at $t=0$. Similar strategies based on multi-valued fixed point theorems have been used in the control literature; see, for instance, the use of Kakutani’s fixed point theorem in [FCZ00].
@@ 6 On the optimality of the regularity
We have the following result.
Claim 6.2.
There is $j_{0}\in\mathbb{N}$ such that for any $j\geq j_{0}$, $\mathbb{P}(A_{j})\geq\kappa\sqrt{|I_{j}|}$ for some positive constant $\kappa$ independent of $j$.
We postpone the proof of this result to the end of this section. Assuming that Claim 6.2 holds, let $A_{N}:=\cap_{j=j_{0}}^{N}A_{j}$ for some natural number $N>j_{0}$. By independence
$\mathbb{P}(A_{N})\geq\kappa^{N}\prod_{j=j_{0}}^{N}\sqrt{|I_{j}|}>0,$ (120)
since it is a finite product. We have that
@@
Moreover, if $\omega\in A_{N}$, then for all $t\in I_{j}$ with $j\in\{j_{0},\ldots,N\}$
@@
$I(\omega)\geq M\quad\text{for all $\omega\in A_{N}$}.$
@@
Note that the probability of the events $A_{N}$ appearing in (120) may become very small as $N$ increases. However, since $N$ is finite, this probability is always strictly positive. This is consistent with the fact that $I<+\infty$ a.s. and highlights that one cannot obtain any uniform bound for the trajectories in $\Omega$.
@@
Proof of Claim 6.2.
By independence of the Brownian increments
$\mathbb{P}(A_{j})=\mathbb{P}\left(\sup_{t\in I_{j}}\left|W_{t}-W_{T-\delta 2^{-j}}\right|\leq\epsilon\sqrt{|I_{j}|}\right)\mathbb{P}\left(|W_{T-\delta 2^{-j}}-c|\leq\epsilon\sqrt{|I_{j}|}\right)=:P_{1}\times P_{2}.$
Let us estimate $P_{1}$. Every $t\in I_{j}$ can be written in the form $t=T-2^{-j}\delta+s2^{-j-1}\delta$ with $s\in[0,1]$. Hence, by the stationarity of the increments of the Brownian increments and that $|I_{j}|=\delta 2^{-j-1}$
$P_{1}=\mathbb{P}\left(\sup_{s\in[0,1]}\left|\frac{W_{s2^{-j-1}\delta}}{\sqrt{\delta 2^{-j-1}}}\right|\leq\epsilon\right).$
Note that $\left\{\frac{W_{s2^{-j-1}\delta}}{\sqrt{\delta 2^{-j-1}}}\right\}\overset{d}{\sim}\left\{\widetilde{W}_{s}\right\}$ for $s\in[0,1]$ where $\widetilde{W}$ is a standard Brownian motion. Therefore,
$P_{1}=\mathbb{P}\left(\sup_{s\in[0,1]}|\widetilde{W}_{s}|\leq\epsilon\right)=\kappa_{\epsilon},$ (124)

### v5

@@ 5.1 The linear case
Note that the randomness in equation (68) appears only through the adapted coefficient $\alpha\in L_{\mathbb{F}}^{\infty}(0,T;L^{\infty}(\mathcal{O}))$ and the $\mathcal{F}_{0}$-measurable initial datum $q_{0}\in L^{p^{\prime}}_{\mathcal{F}_{0}}(\Omega;L^{p^{\prime}}(\mathcal{O}))$. Moreover, for such coefficients, the weak solution $t\mapsto q(t,\cdot)\in L^{2}(\mathcal{O})$ to (68) is $\mathcal{F}_{t}$-measurable for every $t\in[0,T]$.
For $\omega\in\Omega$ fixed, we have from [FZ00, Proposition 3.2]
@@
$\left\lVert q(T)\right\rVert_{L^{p^{\prime}}_{\mathcal{F}_{T}}(\Omega;L^{p^{\prime}}(\mathcal{O}))}\leq C\left\lVert q\right\rVert_{L^{p^{\prime}}_{\mathbb{F}}(0,T;L^{p}(\mathcal{O}_{0}))}.$ (70)
@@
Now, we will follow an argument given by [TZ09], adapted to the $L^{p^{\prime}}$-setting. For any $y_{T}\in L^{p}_{\mathcal{F}_{T}}(\Omega;L^{p}(\mathcal{O}))$, the goal is to find a control $h\in L^{p}_{\mathbb{F}}(0,T;L^{p}(\mathcal{O}_{0}))$ such that the solution $y$ of (67) satisfies $y(0)=0$ a.s.
We introduce the following linear subspace of $L_{\mathbb{F}}^{p^{\prime}}(0,T;L^{p^{\prime}}(\mathcal{O}_{0}))$,
@@
The observability estimate (70) leads to the fact that $\mathcal{L}$ is a bounded linear functional on $\mathcal{X}$ because
$\left|\mathbb{E}\int_{\mathcal{O}}y_{T}q(T)\,\textnormal{d}x\right|\leq\left\lVert y_{T}\right\rVert_{L^{p}_{\mathcal{F}_{T}}(\Omega;L^{p}(\mathcal{O}))}\left\lVert q(T)\right\rVert_{L^{p^{\prime}}_{\mathcal{F}_{T}}(\Omega;L^{p^{\prime}}(\mathcal{O}))}\leq C\left\lVert y_{T}\right\rVert_{L^{p}_{\mathcal{F}_{T}}(\Omega;L^{p}(\mathcal{O}))}\left\lVert q\right\rVert_{L^{p^{\prime}}_{\mathbb{F}}(0,T;L^{p^{\prime}}(\mathcal{O}_{0}))}.$ (71)
By the Hahn-Banach theorem, $\mathcal{L}$ can be extended to a bounded linear functional on $L_{\mathbb{F}}^{p^{\prime}}(0,T;L^{p^{\prime}}(\mathcal{O}_{0}))$ with
$\|\mathcal{L}\|_{\mathcal{L}(L_{\mathbb{F}}^{p^{\prime}}(0,T;L^{p^{\prime}}(\mathcal{O}_{0})),{\mathbb{R}})}\leq C\left\lVert y_{T}\right\rVert_{L^{p}_{\mathcal{F}_{T}}(\Omega;L^{p}(\mathcal{O}))},$ (72)
@@
$(L_{\mathbb{F}}^{p^{\prime}}(0,T;L^{p^{\prime}}(\mathcal{O}_{0}))^{\prime}=L_{\mathbb{F}}^{p}(0,T;L^{p}(\mathcal{O}_{0})).$
This means that there exists $h\in L_{\mathbb{F}}^{p}(0,T;L^{p}(\mathcal{O}_{0}))$ such that
$\mathcal{L}(q)=\mathbb{E}\int_{0}^{T}\int_{\mathcal{O}_{0}}hq\,\textnormal{d}x\textnormal{d}t\qquad\forall q\in L^{p^{\prime}}_{\mathbb{F}}(0,T;L^{p^{\prime}}(\mathcal{O}_{0})),$ (73)
@@
$\mathcal{L}(q)=-\mathbb{E}\int_{\mathcal{O}}y_{T}q(T)\,\textnormal{d}x=\mathbb{E}\int_{0}^{T}\int_{\mathcal{O}_{0}}hq\,\textnormal{d}x\textnormal{d}t\qquad\forall q_{0}\in L^{p^{\prime}}_{\mathcal{F}_{0}}(\Omega;L^{p^{\prime}}(\mathcal{O})).$ (74)
@@
$\mathbb{E}\int_{\mathcal{O}}y(0)q_{0}\,\textnormal{d}x=0\qquad\forall q_{0}\in L^{p^{\prime}}_{\mathcal{F}_{0}}(\Omega;L^{p^{\prime}}(\mathcal{O})).$
@@
$\|h\|_{L_{\mathbb{F}}^{p}(0,T;L^{p}(\mathcal{O}_{0}))}\leq C\left\lVert y_{T}\right\rVert_{L^{p}_{\mathcal{F}_{T}}(\Omega;L^{p}(\mathcal{O}))},$
@@
Now we deal with the more subtle case $p=+\infty$. The difficulty is that one cannot use [LYZ12, Corollary 2.3] because the so-called Radon-Nikodym property is not satisfied by $L^{\infty}$, in particular $(L_{\mathbb{F}}^{1}(0,T;L^{1}(\mathcal{O}_{0}))^{\prime}$ is not equal to $L_{\mathbb{F}}^{\infty}(0,T;L^{\infty}(\mathcal{O}_{0}))$. We start from the $L^{1}$-observability estimate for solution $q$ to (68)
$\left\lVert q(T)\right\rVert_{L^{1}_{\mathbb{F}}(\Omega;L^{1}(\mathcal{O}))}\leq C\left\lVert q\right\rVert_{L^{1}_{\mathbb{F}}(0,T;L^{1}(\mathcal{O}_{0}))},$ (75)
@@
$\|\mathcal{L}\|_{\mathcal{L}(L_{\mathbb{F}}^{1}(0,T;L^{1}(\mathcal{O}_{0})),{\mathbb{R}})}\leq C\left\lVert y_{T}\right\rVert_{L^{\infty}_{\mathcal{F}_{T}}(\Omega;L^{\infty}(\mathcal{O}))},$ (76)
where $C$ is the same as in (75). In particular, $\mathcal{L}$ is a bounded linear functional on $L_{\mathbb{F}}^{2}(0,T;L^{2}(\mathcal{O}_{0}))$ because $L_{\mathbb{F}}^{2}(0,T;L^{2}(\mathcal{O}_{0}))\hookrightarrow L_{\mathbb{F}}^{1}(0,T;L^{1}(\mathcal{O}_{0}))$, therefore we can use the following fact coming from [LYZ12, Corollary 2.3],
$(L_{\mathbb{F}}^{2}(\Omega;L^{2}(0,T;L^{2}(\mathcal{O}_{0})))^{\prime}=L_{\mathbb{F}}^{2}(\Omega;L^{2}(0,T;L^{2}(\mathcal{O}_{0}))).$
This means that there exists $h\in L_{\mathbb{F}}^{2}(\Omega;L^{2}(0,T;L^{2}(\mathcal{O}_{0})))$ such that
$\mathcal{L}(q)=\mathbb{E}\int_{0}^{T}\int_{\mathcal{O}_{0}}hqdxdt\qquad\forall q\in L^{2}_{\mathbb{F}}(0,T;L^{2}(\mathcal{O}_{0})),$ (77)
@@
$\mathcal{L}(q)=-\mathbb{E}\int_{\mathcal{O}}y_{T}q(T)dx=\mathbb{E}\int_{0}^{T}\int_{\mathcal{O}_{0}}hqdxdt\qquad\forall q_{0}\in L^{2}_{\mathcal{F}_{0}}(\Omega;L^{2}(\mathcal{O})).$
@@
Our goal is then to prove that $h\in L_{\mathbb{F}}^{\infty}(0,T;L^{\infty}(\mathcal{O}_{0}))$ and
$\|h\|_{L_{\mathbb{F}}^{\infty}(0,T;L^{\infty}(\mathcal{O}_{0}))}\leq 2C\|y_{T}\|_{L^{\infty}_{\mathcal{F}_{T}}(\Omega;L^{\infty}(\mathcal{O}))}.$ (78)
@@
$\left|\mathbb{E}\int_{0}^{T}\int_{\mathcal{O}_{0}}hqdxdt\right|\leq C\|q\|_{L^{1}_{\mathbb{F}}(0,T;L^{1}(\mathcal{O}))}\|y_{T}\|_{L^{\infty}_{\mathcal{F}_{T}}(\Omega;L^{\infty}(\mathcal{O}))}\quad\forall q\in L^{2}_{\mathbb{F}}(0,T;L^{2}(\mathcal{O}_{0})).$ (79)
@@
$q=\text{sign}(h)\mathbf{1}_{\{|h|>2C\|y_{T}\|_{L^{\infty}_{\mathcal{F}_{T}}(\Omega;L^{\infty}(\mathcal{O}))}\}}\mathbf{1}_{\mathcal{O}_{0}},$
@@
$2C\|y_{T}\|_{L^{\infty}_{\mathcal{F}_{T}}(\Omega;L^{\infty}(\mathcal{O}))}\mu_{\Omega,[0,T],\mathcal{O}_{0}}(|h|>2C\|y_{T}\|_{L^{\infty}_{\mathcal{F}_{T}}(\Omega;L^{\infty}(\mathcal{O}))}),$
@@
$C\|y_{T}\|_{L^{\infty}_{\mathcal{F}_{T}}(\Omega;L^{\infty}(\mathcal{O}))}\mu_{\Omega,[0,T],\mathcal{O}_{0}}(|h|>2C\|y_{T}\|_{L^{\infty}_{\mathcal{F}_{T}}(\Omega;L^{\infty}(\mathcal{O}))}),$
@@
$2C\|y_{T}\|_{L^{\infty}_{\mathcal{F}_{T}}(\Omega;L^{\infty}(\mathcal{O}))}\mu_{\Omega,[0,T],\mathcal{O}_{0}}(|h|>2C\|y_{T}\|_{L^{\infty}_{\mathcal{F}_{T}}(\Omega;L^{\infty}(\mathcal{O}))})\\ \leq C\|y_{T}\|_{L^{\infty}_{\mathcal{F}_{T}}(\Omega;L^{\infty}(\mathcal{O}))}\mu_{\Omega,[0,T],\mathcal{O}_{0}}(|h|>2C\|y_{T}\|_{L^{\infty}_{\mathcal{F}_{T}}(\Omega;L^{\infty}(\mathcal{O}))}).$
@@
$\mu_{\Omega,[0,T],\mathcal{O}_{0}}(|h|>2C\|y_{T}\|_{L^{\infty}_{\mathcal{F}_{T}}(\Omega;L^{\infty}(\mathcal{O}))})=0.$
@@
$|h|\leq 2C\|y_{T}\|_{L^{\infty}_{\mathcal{F}_{T}}(\Omega;L^{\infty}(\mathcal{O}))}\quad\text{a.e. }(t,x)\in Q_{T},\ \text{a.s.}$
@@ 5.2 The nonlinear control problem
The proof of Theorem˜1.11 relies on the so-called source term method, originally developed in [LTT13] for the deterministic setting and later extended to SPDEs in [HLP22]. While [HLP22] already discusses this method for BSPDEs, here we provide a more comprehensive treatment and extend the analysis to our setting.
@@
where $C$ is the constant provided by Proposition˜1.9 (with $p=+\infty$). For some parameters $Q\in(1,\sqrt{2})$ and $R>Q^{2}/(2-Q^{2})$, we define
@@
The proof follows the same steps of [HLP22, Proposition 2.5] with some adaptations to the backward case and the $L^{\infty}$-setting (cf. [LE 20]).
@@
  • In the proof of Theorem˜1.11, we employed a multi-valued fixed point theorem, in contrast to the standard Banach fixed point argument used in Theorem˜1.7. The main reason for this is that Proposition˜5.1 guarantees only the existence of a control that steers the solution of (80) to zero at $t=0$. Similar strategies based on multi-valued fixed point theorems have been used in the control literature; see, for instance, the use of Kakutani’s fixed point theorem in [FZ00].
@@ 6 On the optimality of the regularity
We have the following intermediate result.
Lemma 6.2.
Let $j_{0}\in\mathbb{N}$ be large enough. Define $A^{N}:=\cap_{j=j_{0}}^{N}A_{j}$ for some natural number $N>j_{0}$. Then
$\mathbb{P}(A^{N})>0.$ (119)
We postpone the proof of this result to the end of this section. Assuming that Lemma 6.2 holds, we have that
@@
Moreover, if $\omega\in A^{N}$, then for all $t\in I_{j}$ with $j\in\{j_{0},\ldots,N\}$
@@
$I(\omega)\geq M\quad\text{for all $\omega\in A^{N}$}.$
@@
Note that the probability of the events $A^{N}$ appearing in (119) may become very small as $N$ increases. However, since $N$ is finite, this probability is always strictly positive. This is consistent with the fact that $I<+\infty$ a.s. and highlights that one cannot obtain any uniform bound for the trajectories in $\Omega$.
@@
Proof of Lemma 6.2.
For simplicity, for all $i\in\{j_{0},\ldots,N\}$, we set
$B_{i}$ $:=\left\{\sup_{t\in I_{i}}\left|W_{t}-W_{T-\delta 2^{-i}}\right|\leq\varepsilon\sqrt{|I_{i}|}\right\}$ (123)
therefore $A_{i}:=\left\{\left|W_{T-\delta 2^{-i}}-c\right|\leq\varepsilon\sqrt{|I_{i}|}\right\}\cap B_{i}.$
Let $t_{i}:=T-\delta 2^{-i}$ for $i\in\{j_{0},\ldots,N\}$, and define the function
$\psi_{i}(x):=\mathbb{P}\!\left(\bigcap_{k=i}^{N}A_{k}\,\middle|\,W_{t_{i}}=x\right),\qquad i=j_{0},\ldots,N.$ (124)
The conditioning is understood in the sense of regular conditional probabilities, which are well defined for almost every $x\in\mathbb{R}$ with respect to the law of $W_{t_{i}}$. The function $\psi_{i}$ represents the probability that all remaining events from $i$ to $N$ are satisfied, given the position of the Brownian motion at time $t_{i}$. This interpretation allows us to derive the following recursive representation
$\psi_{i}(x)$ $=\mathbb{E}\!\left(\mathbf{1}_{\bigl\{|W_{t_{i}}-c|\leq\varepsilon\sqrt{|I_{i}|}\bigr\}}\,\mathbf{1}_{B_{i}}\,\mathbf{1}_{\bigcap_{k=i+1}^{N}A_{k}}\;\middle|\;W_{t_{i}}=x\right)$
$=\mathbf{1}_{\bigl\{|x-c|\leq\varepsilon\sqrt{|I_{i}|}\bigr\}}\,\mathbb{E}\!\left(\mathbf{1}_{B_{i}}\,\mathbf{1}_{\bigcap_{k=i+1}^{N}A_{k}}\;\middle|\;W_{t_{i}}=x\right).$ (125)
The identity (6) follows by noting that, under the conditioning $W_{t_{i}}=x$, the constraint on $W_{t_{i}}$ becomes deterministic. Finally, exploiting the tower property of conditional expectation and the Markov property, we get
$\psi_{i}(x)$ $=\mathbf{1}_{\bigl\{|x-c|\leq\varepsilon\sqrt{|I_{i}|}\bigr\}}\mathbb{E}\!\left(\mathbf{1}_{B_{i}}\,\psi_{i+1}(W_{t_{i+1}})\,\middle|\,W_{t_{i}}=x\right),\qquad i=j_{0},\ldots,N-1.$
To obtain the desired result, we look for a strictly positive lower bound for $\psi_{i}$. Observe that if $|x-c|>\varepsilon\sqrt{|I_{i}|}$, then $\psi_{i}(x)=0$. For this reason, it suffices to restrict our attention to $x$ close to $c$ and introduce the interval
$\tilde{I}_{i}:=\left[\,c-\frac{\varepsilon\sqrt{|I_{i}|}}{4},\;c+\frac{\varepsilon\sqrt{|I_{i}|}}{4}\,\right].$ (126)
Observe that if $x\in\tilde{I}_{i}$, then $\mathbf{1}_{\bigl\{|x-c|\leq\varepsilon\sqrt{|I_{i}|}\bigr\}}=1$. We therefore define
$q_{i}:=\inf_{x\in\tilde{I}_{i}}\psi_{i}(x).$ (127)
We now proceed by backward induction to show that the quantities $q_{i}$ are strictly positive for all $i\in\{j_{0},\ldots,N\}$. For the terminal index $i=N$ and for all $x\in\tilde{I}_{N}$, we have $|x-c|\leq\varepsilon\sqrt{|I_{N}|}$. Hence, we note that
$\psi_{N}(x)=\mathbb{P}(B_{N}\mid W_{t_{N}}=x).$
Using the scaling property of Brownian motion we obtain
$\psi_{N}(x)$ $=\mathbb{P}\!\left(\sup_{t\in I_{N}}\left|W_{t}-W_{t_{N}}\right|\leq\varepsilon\sqrt{|I_{N}|}\right)$
$=\mathbb{P}\!\left(\sup_{s\in[0,1]}|\widetilde{W}_{s}|\leq\varepsilon\right)=:\kappa_{\varepsilon},$ (128)
where
$\left\{\frac{W_{t_{N}+s|I_{N}|}-W_{t_{N}}}{\sqrt{|I_{N}|}}\right\}_{s\in[0,1]}\overset{d}{\sim}\{\widetilde{W}_{s}\}_{s\in[0,1]},$